STUDY OF THE ZEEMAN EFFECT IN THE [17.6]7.5 X18.5 TRANSITION IN HOLMIUM MONOXIDE (HoO)

\underline{COLAN LINTON}, Center for Lasers, Atomic and Molecular Sciences and Physics Department, University of New Brunswick, Fredericton, NB Canada E3B 5A3; HAILING WANG AND TIM-OTHY C. STEIMLE , Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287.

The diatomic oxides of the lanthanides and actinides have many low-lying electronic states due to the presence of open f and d orbitals on the metal. The magnetic and electric tuning of the fine structure is an effective means of identifying the dominant configuration of a particular electronic state and of testing theoretical predictions of the configurational composition of the eigenfunctions. Here we report on the first high-resolution molecular beam measurements of the Zeeman effect in electronic transitions of holmium monoxide, HoO. Several branch features in the previously detected [17.6](Ω = 7.5) - X_{1}(Ω = 8.5) electronic transition were recorded at near natural linewidth limit (FWHM 35MHz) field free and in the presence of a tunable static magnetic field. The Zeeman splittings and shifts were used to extract values for the magnitudes of the magnetic g-factors respectively of the two electronic states. These are compared with theoretical predictions.